Find the general solution of the differential equation y^2/2+2ye^t+(y+e^t)dy/dt = 0
1
Expert's answer
2019-03-21T09:51:42-0400

The differential equation can be rewritten as follows:

"\frac12 y^2 + 2 y e^t + \left( y + e^t \right) \frac{dy}{dt} \\ {} = \frac12 y^2 + y e^t + \frac{d}{dt} \left( \frac12 y^2 + y e^t \right) = 0 \, ."

Denoting "u = \frac12 y^2 + y e^t", we have the equation "u + du\/dt = 0", the general solution of which is "u = C_1 e^{-t}", where "C_1" is the integration constant. Hence, we have "\frac12 y^2 + y e^t = C_1 e^{-t}", or "y^2 + 2 y e^t - C e^{-t} = 0", where "C = 2 C_1" is another integration constant. Solving this quadratic equation with respect to y, we obtain the final answer:

"y = - e^t \mp \sqrt{ e^{2t} + C e^{-t}} = - e^t \left( 1 \pm \sqrt{1 + C e^{- 3 t}} \right) \, ."